OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (–2, 4) and is perpendicular to the line y = –2x + 4?

OpenStudy (zale101):

Use the the slope intercept form y = mx + b

OpenStudy (zale101):

m= slope b=intercept

OpenStudy (zale101):

The line y = -2x+4 is in slope-intercept form is -2 = slope 4 = y-intercept

OpenStudy (zale101):

4 = -2(-2) + 4

OpenStudy (anonymous):

first we will identify the slope y = -2x + 4 the slope here is -2. However, you are looking for a perpendicular line, which means the slope is going to be the negative reciprocal of that slope....to get the negative reciprocal, all you do is " flip " the slope and change the sign. Therefore, the slope of the perpendicular line is negative reciprocal of -2, which is 1/2. Now we will use the point slope form because we have the points and the slope. (-2,4) x1 = -2, y1 = 4 slope(m) = 1/2 y - y1 = m(x - x1) now lets sub our info into the formula.. y - 4 = 1/2(x - (-2) y - 4 = 1/2(x + 2) y - 4 = 1/2x + 1 y = 1/2x + 1 + 4 y = 1/2x + 5 <-- this is your equation